Jennaofcbaby:

A triangle has a base of 2^3 units and a height of 2^4 units. What is the area of the​ triangle? Write your answer as a power of 2

supie:

Do you know the formula for the area of a triangle?

Jennaofcbaby:

here you go the formula

Jennaofcbaby:

@supie

supie:

So, we have to do \(Base*Hight\) Correct?

supie:

@jimthompson5910

Jennaofcbaby:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So, we have to do \(Base*Hight\) Correct? \(\color{#0cbb34}{\text{End of Quote}}\) I'm pretty sure

Jennaofcbaby:

who is that you just @?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Jennaofcbaby who is that you just @? \(\color{#0cbb34}{\text{End of Quote}}\) A smart guy.

Jennaofcbaby:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @Jennaofcbaby who is that you just @? \(\color{#0cbb34}{\text{End of Quote}}\) A smart guy. \(\color{#0cbb34}{\text{End of Quote}}\) Oh cool.

Jennaofcbaby:

Oop imma get an A with him.

jimthompson5910:

You're close, but the area of a triangle is \(\large \frac{\text{base*height}}{2}\)

Jennaofcbaby:

Okay but i need it written in power of 2

Jennaofcbaby:

If that's it bc im not good at math

jimthompson5910:

If base = 2^3 and height = 2^4, then we can say: \(\large \text{area} = \frac{\text{base*height}}{2}\) \(\large \text{area} = \frac{2^3*2^4}{2}\) \(\large \text{area} = \frac{2^3*2^4}{2^1}\) What's the next step from here?

Jennaofcbaby:

to be honest I really don't know but what it says it's just I have to write it in the power to

Jennaofcbaby:

@jimthompson5910

jimthompson5910:

Recall that we have these exponent rules \(\large a^b*a^c = a^{b+c}\) and \(\large \frac{a^b}{a^c} = a^{b-c}\) which I'll refer to as rule 1 and rule 2 respectively Using those rules leads to the following: \(\large \text{area} = \frac{2^3*2^4}{2^1}\) \(\large \text{area} = \frac{2^{3+4}}{2^1} .... \text{ use rule 1}\) \(\large \text{area} = \frac{2^7}{2^1}\) \(\large \text{area} = 2^{7-1} .... \text{ use rule 2}\) There's one more step after this.

Jennaofcbaby:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 Recall that we have these exponent rules \(\large a^b*a^c = a^{b+c}\) and \(\large \frac{a^b}{a^c} = a^{b-c}\) which I'll refer to as rule 1 and rule 2 respectively Using those rules leads to the following: \(\large \text{area} = \frac{2^3*2^4}{2^1}\) \(\large \text{area} = \frac{2^{3+4}}{2^1} .... \text{ use rule 1}\) \(\large \text{area} = \frac{2^7}{2^1}\) \(\large \text{area} = 2^{7-1} .... \text{ use rule 2}\) There's one more step after this. \(\color{#0cbb34}{\text{End of Quote}}\) Look at the ss i sent that's what i see

jimthompson5910:

I've looked at it. You just need to simplify \(\large 2^{7-1}\) at this point.

Jennaofcbaby:

So what is it?

jimthompson5910:

Simplifying the 7-1 leads to 6, so \(\large 2^{7-1} = 2^6\)

Jennaofcbaby:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 Simplifying the 7-1 leads to 6, so \(\large 2^{7-1} = 2^6\) \(\color{#0cbb34}{\text{End of Quote}}\) In what unit? square units cubic units or just units

jimthompson5910:

areas always deal with square units eg: square cm, square inches, etc

Jennaofcbaby:

Okay thanks for your help!

jimthompson5910:

You're welcome